OpenStudy (anonymous):

Find the derivative of the function: f(x)= ln(1/7x^2) I'm figuring that i would subtract the top of the fraction by the bottom, but then I get stuck as to how to get the derivative.. so confusing

OpenStudy (anonymous):

2/x

OpenStudy (anonymous):

\[\large f(x)= \ln(\frac {1}{7x^2})\] is this your function?

OpenStudy (anonymous):

@dpaInc yep

sam (.sam.):

wait its 7x^2 lol

OpenStudy (anonymous):

f(x) = ln1 - ln(7x^2) f'(x) = -[7x^2]'/[7x^2] = -14x/(7x^2) = -2/x

OpenStudy (anonymous):

so ln1 = 0??

OpenStudy (anonymous):

and i just take the derivative of the second part?

sam (.sam.):

yep

sam (.sam.):

-2/x?

OpenStudy (anonymous):

yeah my online hw took that answer as correct. just trying to understand the how for my exam..

OpenStudy (anonymous):

i get the derivatives rules, but when "ln" is thrown in the mix, i get lost

OpenStudy (anonymous):

should i treat ln as a type of variable? I can't wrap my head around natural log very wel